# The open driven two-level system at conical intersections of quasienergies

### v1

@@ Abstract
We study the stationary state of an ac-driven two-level system under particle exchange with a fermionic environment. A particular question addressed is whether there exist limits in which the populations of the Floquet states are determined by their quasienergies or their mean energies, respectively. The focus lies on parameters in the vicinity of conical intersections of quasienergies, because there the two kinds of energies behave rather differently, such that the characteristics of the two intuitive limits are most pronounced. A main finding is a crossover from a Floquet-Gibbs state at low temperatures to a mean-energy dominated state at intermediate temperatures. Analytical estimates are confirmed by numerical calculations.
@@ I Introduction
Quantum systems at thermal equilibrium are usually assumed to be in a Gibbs states, depending on whether one allows for particle exchange, it will be the canonical or the grand canonical ensemble. For a more profound description, one may employ system-bath models that allow for the exchange of energy Weiss (2012) or of particles Bruus and Flensberg (2004) with an environment to obtain the dynamics and the stationary solution of the reduced system density operator. Usually at least for very weak coupling, the latter eventually becomes a Gibbs state.Landi, Poletti, and Schaller (2022) For driven systems, the situation is more involved, because no such formal expression for the stationary density operator is known. Hence, for its computation, system-bath models are indispensable. Notably, then even in the weak-coupling limit, the stationary state may qualitatively depend on details of the system-bath interaction.Ferrón, Domínguez, and Sánchez (2012); Blattmann, Hänggi, and Kohler (2015) Moreover, genuine non-equilibrium effects such as pumping and rectification of heat Segal and Nitzan (2005, 2006) and charge Lehmann et al. (2002, 2003) emerge. Such non-equilibrium effects are particularly important for Floquet engineering which is the design of effective static Hamiltonians by ac fields.Grossmann et al. (1991); Großmann and Hänggi (1992); Holthaus (1992); Eckardt (2017) In exceptional cases of driven dissipative systems, there emerge Floquet-Gibbs states, i.e., canonical states in which the eigenenergies are formally replaced by quasienergies.Thingna, Wang, and Hänggi (2012); Shirai, Mori, and Miyashita (2015); Shirai et al. (2016); Iwahori and Kawakami (2016); Mori (2023) There are also situations in which, by contrast, the mean energies of the Floquet states determine the populations.Kohler et al. (1998); Ketzmerick and Wustmann (2010); Kohler (2024)
In practice none of these limits will be perfectly realized. Nevertheless one may observe a tendency towards the one or the other. To explore this behavior, the vicinity of conical intersections of quasienergies turned out rather to be interesting.Kohler (2024) Such intersections emerge when a spatio-temporal symmetry of a driven quantum system such as generalized parity Peres (1991) allows the exact crossing of quasienergies. A weak perturbation may break this symmetry, so that the former crossing turns into an accidental degeneracy in a two-dimensional parameter space. A most interesting feature of canonical intersections in driven systems is that along lines with constant quasienergies, the mean energies of the Floquet states are interchanged.Kohler (2024) Therefore in this regime, the roles played by these energies will be most pronounced. For example, there may emerge discontinuities of the populations Engelhardt, Platero, and Cao (2019) or a crossover from a mean-energy state to a more exceptional Floquet-Gibbs state.Kohler (2024)
@@
This work is structured as follows. In Sec. II, we introduce conical intersections of quasienergies as a consequence of a spatio-temporal symmetry, the coupling to a fermionic particle reservoir, and a master equation formalism for the single-particle density operator based on Floquet theory. Section III is devoted to approximative solutions that lead to conjectures for the stationary state which are confirmed by numerical calculations. Conclusions are drawn in Sec. IV, while the properties of Floquet states close to conical intersections are summarized in the appendix.
@@ II.1 Driven two-level system and conical intersections
A suitable tool for treating periodically time-dependent quantum systems is Floquet theory. Its cornerstone is discrete time translation invariance, from which follows that the Schrödinger equation has a complete set of solutions of the form $e^{-iqt}|\phi(t)\rangle$ with quasienergy $q$ and Floquet state $|\phi(t)\rangle=|\phi(t+2\pi/\Omega)\rangle$. Shirley (1965); Sambe (1973); Hänggi (1998) The Floquet states and their quasienergies are eigensolutions of the operator $\mathcal{H}=H(t)-i\partial/\partial t$ in Sambe space, which is the direct product of Hilbert space and the space of $2\pi/\Omega$-periodic functions.Sambe (1973) Quasienergies correspond to the phase that a Floquet state acquires during one driving period. An important property of a Floquet state is its mean energy $E$, i.e, its energy expectation value averaged over one driving period. It relates to the quasienergy via a geometric phase,Aharonov and Anandan (1987); Moore (1990)
@@
where the outer angles denote time average, while $|\phi_{k}\rangle$ is the $k$th Fourier coefficient of the Floquet state.
@@ II.2 Fermionic environment and master equation
For very weak coupling one can apply a rotating-wave approximation which ignores off-diagonal elements of $R$ such that $R_{\alpha\beta}=P_{\alpha}\delta_{\alpha,\beta}$ with the Floquet state populations $P_{\alpha}$. Then in the long-time limit, the populations become practically time-independent and obey Lehmann et al. (2003)
@@ III.1 Equal coupling to both leads
Interestingly, the population is independent of $\Gamma$, despite that the wire-lead couplings $V_{\ell}$ do not commute with the wire Hamiltonian. This result is consistent with the natural expectation for the undriven limit in which the occupation probability of the sidebands becomes $\propto\delta_{k,0}$. Then $P_{\alpha}=f(q_{\alpha})$ with the quasienergy $q_{\alpha}$ being the corresponding eigenenergy of the undriven Hamiltonian.
@@
To demonstrate that this discrepancy can be attributed to the relatively small driving frequency, we compare the conjectures (12) and (13) with the numerical results for different driving frequencies and temperatures. As quantity of interest, we consider the population difference $\Delta P\equiv|P_{1}-P_{0}|$ for an amplitude slightly above the cone tip, i.e., for $\vartheta=\pi$. The results expected from Eqs. (12) and (13) are $\Delta P=\langle\phi_{\alpha 0}|\phi_{\alpha 0}\rangle\coth(|q_{0}|/2k_{B}T)$ and $\coth(|E_{0}|/2k_{B}T)$, respectively, where we have used $q_{0}+q_{1}=0=E_{0}+E_{1}$. Figure 4 shows that the approximation in Eq. (13) is fulfilled rather well for $k_{B}T\gtrsim\Omega$. However, the Floquet-Gibbs limit (dotted line) only emerges for a relatively large driving frequency $\Omega\ggg\Delta$. In this respect, the present situation is analogous to the one with a bosonic heat bath.Kohler (2024) It is worth mentioning that same conclusions can be drawn from the data for $\vartheta=0$. For this value, however, quasienergies and mean energies have opposite order, $q_{0}<q_{1}$ while $E_{0}>E_{1}$. This leads to a non-monotonic behavior as a function of $k_{B}T$ which is less convenient to analyze.
@@ III.2 Asymmetric wire-lead couplings
To highlight the role of equal coupling to both leads, we consider the asymmetric case with the right lead disconnected, $\Gamma_{R}=0$. All other parameters are as before, such that both the quasienergies and the mean energies are unchanged. Nevertheless, the resulting Floquet state populations shown in Fig. 5(a) differ qualitatively from the ones in Fig. 3(b). Even the reflection symmetry at $\vartheta=\pi$ is lost. Obviously, we nor longer have a Floquet-Gibbs-like state or a mean energy state.
@@
Figure  5(b) depicts one of the prefactors of the Fermi function in Eq. (14) for the sideband indices $k=0,\pm 1$. It can be appreciated that at $\vartheta=\pi/2$, they are rather small and for some value of $\vartheta$ practically vanish. However, they vanish at different angles. Therefore, their relative difference may be huge, and the symmetry with respect to the sign of $k$ underlying the derivation of Eq. (12) may be significantly broken. As a consequence, a dependence of $P_{\alpha}$ on some sideband energies will remain.
For the high-temperature approximation that leads to the mean-energy state the situation is even clearer. The sideband occupation in the middle of Eq. (13) now contains the projector $|L\rangle\langle L|$. Therefore, the summation no longer yields the mean energy. This is already gradually the case for $\Gamma_{L}\neq\Gamma_{R}$. This reasoning reveals that the symmetry of the dot-lead couplings is a crucial ingredient for a stationary state that can be written as simple function of the quasienergy or the mean energy.
Let finally us remark that the (approximate) localization of the Floquet states for $A=A_{0}$ and small but non-vanishing detuning $\epsilon$ (i.e., $\vartheta=\pi/2$ and $23\pi/$ 2) that explains the data in Fig. 5(b) can be understood form the perturbation theory in Sec. II.1. Since at the tip, the zeroth order of the Floquet Hamiltonian vanishes, the Floquet states are essentially determined by the detuning, which is proportional to $\sigma_{z}$. Hence the natural expectation is a tendency towards localized states.
@@ IV Conclusions
We have addressed the question whether for a driven quantum system coupled to electron reservoirs, Floquet-Gibbs states or mean-energy states emerge. As a basic model, we have considered a two-level system that can be occupied with up to two spinless electrons without taking their interaction into account. This situation allows a good comparison with a former study Kohler (2024) in which dissipation stems from a coupling to a bosonic heat bath. Also here we have focussed on conical intersections, because in their vicinity, quasienergies and mean energies behave rather differently and in a characteristic manner. Then much insight can be gained from the Bloch-Redfield master equation for the single-particle density operator. Within a rotating-wave approximation it can be solved analytically.
A main observation is that no true Floquet-Gibbs states are found, i.e., there seem to be no non-trivial limit in which the populations match the Fermi function evaluated at the quasienergies. Nevertheless, there are situations in which the deviation is determined by the quasienergy in the first Brillouin zone, such that one may speak of Floquet-Gibbs-like states. As for the dissipative two-level system, this requires rather large driving frequencies. For smaller frequencies and intermediate temperatures, the stationary populations are directly given by a Fermi function with the mean energy of the Floquet states as argument. Thus, also here the mean-energy states seem rather generic, at least under the condition of a sufficiently symmetric setup. In particular, the tunnel couplings of the sites to the respective leads have to be equal. For larger systems, for example for models that describe conducting molecules, such equal coupling of all sites to an electron reservoir may be essential for the emergence of Floquet-Gibbs-like states or mean-energy states.
@@ Appendix A Floquet Hamiltonian near the cone tip
with $A_{0}$ assuming a value for which the quasienergies with $q=0$ are degenerate. Since $\mathcal{H}_{0}$ obeys generalized parity, $G\mathcal{H}_{0}G^{-1}=\mathcal{H}_{0}$, its two non-equivalent Floquet states can be classified as even or odd. They will be denoted by $|\varphi_{\pm}(t)\rangle$ with mean energies $E_{\pm}$. The perturbation

### v2

@@ Abstract
We study the stationary state of an ac-driven two-level system under particle exchange with a fermionic environment. A particular question addressed is whether there exist limits in which the populations of the Floquet states are determined by their quasienergies or their mean energies, respectively. The focus lies on parameters in the vicinity of conical intersections of quasienergies, because there the two kinds of energies behave rather differently, such that the characteristics of the two intuitive limits are most pronounced. A main finding is a crossover from a Floquet-Gibbs-like state at low temperatures to a mean-energy dominated state at intermediate temperatures. Analytical estimates are confirmed by numerical calculations.
@@ I Introduction
Quantum systems at thermal equilibrium are usually assumed to be in a Gibbs state which, depending on whether one allows for particle exchange, is the canonical or the grand canonical ensemble. For a more profound description, one may employ system-bath models that allow for the exchange of energy Weiss (2012) or particles Bruus and Flensberg (2004) with an environment to obtain the dynamics and the stationary solution of the reduced system density operator. Usually at least for very weak coupling, the latter eventually becomes a Gibbs state.Landi, Poletti, and Schaller (2022) For driven systems, the situation is more involved, because no such formal expression for the stationary density operator is known. Hence, for its computation, system-bath models are indispensable. Notably, then even in the weak-coupling limit, the stationary state may qualitatively depend on details of the system-bath interaction.Ferrón, Domínguez, and Sánchez (2012); Blattmann, Hänggi, and Kohler (2015) Moreover, genuine non-equilibrium effects emerge such as pumping and rectification of heat Segal and Nitzan (2005, 2006) and charge Lehmann et al. (2002, 2003) currents. These non-equilibrium effects are particularly important for Floquet engineering which is the design of effective static Hamiltonians by ac fields.Grossmann et al. (1991); Großmann and Hänggi (1992); Holthaus (1992); Eckardt (2017) In exceptional cases, driven dissipative systems assume Floquet-Gibbs states, i.e., canonical states in which the eigenenergies are formally replaced by quasienergies.Thingna, Wang, and Hänggi (2012); Shirai, Mori, and Miyashita (2015); Shirai et al. (2016); Iwahori and Kawakami (2016); Mori (2023) There are also situations in which, by contrast, the mean energies of the Floquet states determine the populations.Kohler et al. (1998); Ketzmerick and Wustmann (2010); Kohler (2024)
In practice none of these limits will be perfectly realized. Nevertheless one may observe a clear tendency towards the one or the other. To explore this behavior, the vicinity of conical intersections of quasienergies turned out to be rather interesting.Kohler (2024) Such intersections emerge when a spatio-temporal symmetry of a driven quantum system such as generalized parity Peres (1991) allows the exact crossing of quasienergies. A weak perturbation may break this symmetry, so that the former crossing turns into an accidental degeneracy in a two-dimensional parameter space. A most interesting feature of canonical intersections in driven systems is that along lines with constant quasienergies, the mean energies of the Floquet states are interchanged.Kohler (2024) Therefore in this regime, the roles played by these energies will be most pronounced. For example, there may emerge discontinuities of the populations Engelhardt, Platero, and Cao (2019) or a crossover from a mean-energy state to a more exceptional Floquet-Gibbs state.Kohler (2024)
@@
This work is structured as follows. In Sec. II, we introduce conical intersections of quasienergies as a consequence of a spatio-temporal symmetry, the coupling to a fermionic particle reservoir, and a master equation formalism for the single-particle density operator based on Floquet theory. Section III is devoted to approximative solutions of the master equation which lead to conjectures for the stationary state which are confirmed by numerical calculations. Conclusions are drawn in Sec. IV, while the properties of Floquet states close to conical intersections are summarized in the appendix.
@@ II.1 Driven two-level system and conical intersections
A suitable tool for treating periodically time-dependent quantum systems is Floquet theory. Its cornerstone is discrete time translation invariance from which follows that the Schrödinger equation has a complete set of solutions of the form $e^{-iqt}|\phi(t)\rangle$ with quasienergy $q$ and Floquet state $|\phi(t)\rangle=|\phi(t+2\pi/\Omega)\rangle$. Shirley (1965); Sambe (1973); Hänggi (1998) The Floquet states and their quasienergies are eigensolutions of the operator $\mathcal{H}=H(t)-i\partial/\partial t$ in Sambe space, which is the direct product of Hilbert space and the space of $2\pi/\Omega$-periodic functions.Sambe (1973) Quasienergies correspond to the phase that a Floquet state acquires during one driving period. An important property of a Floquet state is its mean energy $E$, i.e, its energy expectation value averaged over one driving period. It relates to the quasienergy via a geometric phase,Aharonov and Anandan (1987); Moore (1990)
@@
where the outer angles denote time average, while $|\phi_{k}\rangle$ is the $k$th Fourier coefficient of the Floquet state $|\phi(t)\rangle$.
@@ II.2 Fermionic environment and master equation
For very weak coupling one can apply a rotating-wave approximation which ignores off-diagonal elements of $R$ such that $R_{\alpha\beta}=P_{\alpha}\delta_{\alpha,\beta}$ with the Floquet state populations $P_{\alpha}$. In the long-time limit, the populations become practically time-independent and obey Lehmann et al. (2003)
@@
An alternative derivation of Eq. (8) starts from a formulation with Green’s functions for which the Floquet equation of the two-level system contains a self-energy stemming from the coupling to the leads. Kohler, Lehmann, and Hänggi (2005); Arrachea and Moskalets (2006); Stefanucci et al. (2008) It allows a treatment beyond weak site-lead coupling and, thus, considers the resulting level broadening. It has been shown that for weak coupling, the Floquet-Green’s function approach becomes equivalent to the present master equation. Kohler, Lehmann, and Hänggi (2005) Notice that strong coupling is not considered here, because already without the driving, it causes significant deviations from a Gibbs state.Chung and Peschel (2001); Cheong and Henley (2004); Sharma and Rabani (2015)
@@ III.1 Equal coupling to both leads
Interestingly, the populations are independent of $\Gamma$, despite that the wire-lead couplings $V_{\ell}$ do not commute with the wire Hamiltonian. This result is consistent with the natural expectation for the undriven limit in which the occupation probability of the sidebands becomes $\propto\delta_{k,0}$. Then $P_{\alpha}=f(q_{\alpha})$ with the quasienergy $q_{\alpha}$ being the corresponding eigenenergy of the undriven Hamiltonian.
@@
To demonstrate that this discrepancy can be attributed to the relatively small driving frequency, we compare the conjectures (12) and (13) with the numerical results for different driving frequencies and temperatures. As quantity of interest, we consider the population difference $\Delta P\equiv|P_{1}-P_{0}|$ for an amplitude slightly above the cone tip, i.e., for $\vartheta=\pi$. The results expected from Eqs. (12) and (13) are $\Delta P=\langle\phi_{\alpha 0}|\phi_{\alpha 0}\rangle\coth(|q_{0}|/2k_{B}T)$ and $\coth(|E_{0}|/2k_{B}T)$, respectively, where we have used $q_{0}+q_{1}=0=E_{0}+E_{1}$. Figure 4 shows that the approximation in Eq. (13) is fulfilled rather well for $k_{B}T\gtrsim\Omega$. However, the Floquet-Gibbs limit (dotted line) only emerges for a relatively large driving frequency $\Omega\ggg\Delta$. In this respect, the present situation is analogous to the one with a bosonic heat bath.Kohler (2024) The same conclusions can be drawn from the data for $\vartheta=0$. For this value, however, quasienergies and mean energies have opposite order, $q_{0}<q_{1}$ while $E_{0}>E_{1}$. This leads to a non-monotonic behavior as a function of $k_{B}T$ which is less convenient to analyze.
@@ III.2 Asymmetric wire-lead couplings
To highlight the role of equal coupling to both leads, we consider the asymmetric case with the right lead disconnected, $\Gamma_{R}=0$. All other parameters are as before, such that both the quasienergies and the mean energies are unchanged. Nevertheless, the resulting Floquet state populations shown in Fig. 5(a) differ qualitatively from the ones in Fig. 3(b). Even the reflection symmetry at $\vartheta=\pi$ is lost. Obviously, we no longer obtain a Floquet-Gibbs-like state or a mean energy state.
@@
Figure  5(b) depicts one of the prefactors of the Fermi function in Eq. (14) for the sideband indices $k=0,\pm 1$. It can be appreciated that close to $\vartheta=\pi/2$, they are rather small and for some value of $\vartheta$ practically vanish. However, they vanish at different angles. Therefore, their relative difference may be huge, and the symmetry with respect to the sign of $k$ underlying the derivation of Eq. (12) may be significantly broken. As a consequence, a dependence of $P_{\alpha}$ on some sideband energies will remain.
For the high-temperature approximation which leads to the mean-energy state the situation is even clearer. The sideband occupation in the middle of Eq. (13) now contains the projector $|L\rangle\langle L|$. Therefore, the summation no longer yields the mean energy. This is already gradually the case for $\Gamma_{L}\neq\Gamma_{R}$. This reasoning reveals that the symmetry of the dot-lead couplings is a crucial ingredient for a stationary state that can be written as simple function of the quasienergy or the mean energy.
Let us finally remark that the (approximate) localization of the Floquet states for $A=A_{0}$ and small but non-vanishing detuning $\epsilon$ (i.e., $\vartheta=\pi/2$ and $23\pi/$ 2) that explains the data in Fig. 5(b) can be understood form the perturbation theory in Sec. II.1. Since at the tip, the zeroth order of the Floquet Hamiltonian vanishes, the Floquet states are essentially determined by the detuning, which is proportional to $\sigma_{z}$. Hence the natural expectation is a tendency towards localized states.
@@ IV Conclusions
We have addressed the question whether for a driven quantum system coupled to electron reservoirs, Floquet-Gibbs states or mean-energy states emerge. As a basic model, we have employed a two-level system that can be occupied with up to two spinless electrons without taking their interaction into account. This situation allows a good comparison with a former study Kohler (2024) in which dissipation stems from a coupling to a bosonic heat bath. Also here we have focussed on conical intersections, because in their vicinity, quasienergies and mean energies behave rather differently and in a characteristic manner. Then much insight can be gained from the Bloch-Redfield master equation for the single-particle density operator. Within a rotating-wave approximation it has been solved analytically.
A main observation is that no true Floquet-Gibbs states are found, i.e., the populations match the Fermi function evaluated at the quasienergies only in trivial limits such as in the absence of driving. Nevertheless, there are situations in which the deviation is determined by the quasienergy in the first Brillouin zone, such that one may speak of Floquet-Gibbs-like states. As for the dissipative two-level system, this requires rather large driving frequencies. For smaller frequencies and intermediate temperatures, the stationary populations are directly given by a Fermi function with the mean energy of the Floquet states as argument. Thus, also here the mean-energy states seem rather generic, at least under the condition of a sufficiently symmetric setup. In particular, the tunnel couplings of the sites to the respective leads have to be approximately equal. For larger systems, for example for models that describe conducting molecules, such equal coupling of all sites to an electron reservoir may be essential for the emergence of Floquet-Gibbs-like states or mean-energy states.
Our prediction of certain reduced density operators may be tested in any experiment that is sensitive to the occupation of Floquet states. One may for example proceed as in Ref. Chen et al., 2021 where a driven closed double quantum dot occupied with a single electron was coupled to a superconducting cavity. Then the transmission of the latter provides information about the occupation of Floquet states. A corresponding experiment with an open double dot may provide novel insight to the stationary state of driven quantum systems.
@@ Appendix A Floquet Hamiltonian near the cone tip
with $A_{0}$ such that the quasienergies cross, i.e., they are degenerate and vanish, $q_{0}=q_{1}=0$. Since $\mathcal{H}_{0}$ obeys generalized parity, $G\mathcal{H}_{0}G^{-1}=\mathcal{H}_{0}$, its two non-equivalent Floquet states can be classified as even or odd. They will be denoted by $|\varphi_{\pm}(t)\rangle$ and their mean energies bay $E_{\pm}$. The perturbation